OpenStudy (anonymous):

If f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

ineed an aswer please

OpenStudy (frostbite):

Why not post this in the math section?

OpenStudy (frostbite):

Anyway: f(g(x)= 2(8−x) − 2(8−x) + 9 = 16-2x=9 f(g(-4))=9 if my calculations are right

OpenStudy (frostbite):

eh forgot f(x) can be rewritten to 0+9.. therefore the answer is ofcuase always 9

OpenStudy (anonymous):

How did I get 111?

OpenStudy (anonymous):

Becasuse if x = -4, then the first thing we need to do is solve for g(x), right? g(x) is 8 - x. So, if x = -4, then g(x) = 12. Right so far?

OpenStudy (anonymous):

And THEN we take 12, and we plug it in for all the X's, in f(x).

OpenStudy (anonymous):

That's 12^2 - 2(12) + 9

OpenStudy (anonymous):

We add before we subtract. So that'd be... 144 - 33. And that's 111.

OpenStudy (frostbite):

oh is it x^2.... i thought it was 2 times x

OpenStudy (frostbite):

2x x2 acfsd we see as multuiplication in mathmatics. 2*x, x*2, a*c*f*s*d.

OpenStudy (anonymous):

Yeah I know what you mean. I also knew that whenever we multiply X with 2, it's usually shown as 2x.

OpenStudy (frostbite):

Hmmm then you need quite alot of good will.

OpenStudy (anonymous):

@bobbya Did you mean to say "X squared," when you wrote "x2"?

OpenStudy (anonymous):

@Frostbite Good will?

OpenStudy (frostbite):

I just ment i calculated what I saw :P

OpenStudy (frostbite):

Not what I had to guess he ment ^^

OpenStudy (anonymous):

I get it. Just curious really.